# suppReferences

### v1

@@
Recent developments in structured light have enabled the generation of optical Skyrmions [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15], which are complex spatially varying polarization fields that carry information through a topological number taking values in the integers [16]. Optical Skyrmions have three crucial properties that make them ideal candidates for high-density data transfer [17, 16, 15], namely the ability to interface with digital information given the discrete nature of the Skyrmion number, a general robustness to perturbations, and the potential to store arbitrarily large integers within a single localized analog optical field.
@@
These same reasons also make the optical Skyrmion a natural candidate for computation, with its integer-valued topological number offering a route to digital photonic computing in a way that transcends the usual bitwise framework of digital electronics. Moreover, compared to existing photonic computing strategies which predominantly modulate amplitude and phase, complex spatially varying polarization fields represent an untapped dimension that can be independently manipulated, and therefore has the potential to increase bandwidth significantly.
@@
The central reason for exploring the use of optical Skyrmions in photonic computing is, however, their intrinsic resistance to noise and perturbations which arises from the integer-valued nature of the Skyrmion number, with topological robustness already established in various settings [18, 19], including free-space propagation in the paraxial limit [20, 21], and through various complex media such as spatially varying retarders, diattenuators, and depolarizers [16]. This robustness to noise is particularly relevant to photonic computing, which has thus far remained analog and therefore susceptible to noise. Take, for instance, the following three common photonic accelerators. The accumulation of random phase errors in Mach-Zehnder interferometer meshes [22, 23], thermal crosstalk in microring resonator weight banks [24, 25] and low optical contrast in phase-change material-based photonic crossbar arrays [26] all lead to degraded signal-to-noise ratio in large-size implementations, circumventing the scalability of these architectures. Even with progress in device innovation [27, 28], system calibration, and control algorithm optimization to minimize errors[29, 30], almost all demonstrated photonic accelerators thus far have had a size of less than 4 inputs by 4 outputs. Computing using optical Skyrmions is therefore a promising solution offering greater scalability through improved stability against noise.
@@
Motivated by this possibility, we describe here a class of structured matter which when restricted to particular types of Skyrmions behave as adders/subtractors (Fig. 1) and provide experimental evidence supporting this fact. Note that the same medium can perform both addition and subtraction, but we adopt the term “adder” throughout this paper for brevity. Such structured matter can be realized using spatially varying retarders with specific structures on their boundary (Extended Fig. 1), and offer a reliable way of manipulating the Skyrmion number without external energy input. Our work opens the doors to this entirely new framework for performing photonic computing, where optical Skyrmions are used as the fundamental unit in computation.
@@
Fig. 1: Concept. An optical Skyrmion of arbitrary order $n$ passing through specially designed passive structured matter can have the effect of addition or subtraction by an arbitrary integer, $n\mapsto n\pm k$. Note that the same medium can be used to perform both addition and subtraction. Such structured matter can be realized in many different ways including metasurfaces, inkjet printing of liquid crystal structures, gradient index systems, direct laser writing of birefringent structures in silica, and more. This figure also depicts the Stokes fields of a standard Bloch-type Skyrmion passing through such a medium. Throughout this paper, color is used to represent azimuthal angle on the Poincaré sphere, and saturation to represent height.
@@ 1 Main
Note our mathematical results also imply that the function of any medium designed in this way depends only on the structure of its boundary and is independent of material properties everywhere else provided they are continuous. This is a reflection of the topological structure of the optical Skyrmion exhibited in matter (see details in Discussion and Methods). From a practical perspective, this suggests a strong robustness of any such adder to physical imperfections of the medium that implements it, which greatly eases fabrication.
To give a concrete example, consider a continuously varying retarder [31], which when restricted to its boundary, is linear and has a retardance of $\pi$. Let $k$ be the number of half-revolutions made by the axis of the retarder traversing counterclockwise along its boundary. Then one may show
@@
To cascade multiple adders together, one must realign the outer boundary after each operation which is most easily done using a homogenous waveplate. With our proposed design, this amounts to adding a half waveplate after the medium for addition and a half waveplate before for subtraction, provided we design our components for Skyrmions that are right circularly polarized on their boundary (see Extended Fig. 2a). Therefore, combining an adder with a half waveplate effectively produces a cascadable module that performs addition in one direction and subtraction in the other.
Lastly, to demonstrate the feasibility of our proposed medium, we provide experimentally measured Stokes fields of Skyrmions passing through second-order adders in Fig. 2. We generated Skyrmions of orders ranging from $-3$ to 3 using a cascade of 2 spatial light modulators [10] and passed the field through the medium with appropriate waveplates to achieve the operations $+2+2$, $+2-2$, $-2+2$ and $-2-2$ (see Extended Fig. 2b). Note the medium we used was discarded by quality assurance due to its asymmetric axis distribution, and it exhibits significant perturbations in material properties compared to the typical sample. Despite this, the numerically computed Skyrmion numbers in Fig. 2 show that our proposed adder efficiently and reliably performs the desired operations, even with imperfections in the medium.
@@ 2 Discussion
As mentioned earlier, a crucial property of any such adder is that its effect on Skyrmion fields depends only on the structure of its boundary, a reflection of the topological nature of the optical Skyrmion. From an engineering perspective, this provides clear advantages in manufacturing as the tolerable margins of error in fabrication are greatly relaxed, with no further restrictions on the medium besides the continuity of its material properties. In particular, the function of the adder is independent of perturbations to material properties away from the boundary, including any spatially varying anisotropic absorptions and complex spatially varying retardance such as those resulting from birefringence [16]. We demonstrate the robustness arising from this “topological duality” between field and matter in Extended Fig. 1, which shows the invariance of the effect of our proposed adder on the Skyrmion number of optical fields to disorder in the medium.
@@
Secondly, understanding the non-trivial behavior of polarization fields in waveguides is of key importance for enabling small integrated devices to make use of Skyrmions. In particular, apart from fabricating the adders themselves, it is also necessary to relay polarization information between adders for any meaningful computation to be done. For a waveguide to support the propagation of complex structured fields, it must necessarily be large enough to carry multiple modes. Moreover, work will need to be done to establish the range of Skyrmion numbers that can be achieved by an arbitrary superposition of propagating modes, and the conservation of Skyrmion number in propagation. Here, we provide a heuristic argument suggesting the feasibility of topological protection. Based on the ellipticity of the Helmholtz equation, one has in general that the electric field develops continuously within the waveguide. This then naturally descends onto a homotopy of Stokes fields provided there are no zeros of the field. Therefore, at least for some finite distance, we expect the Skyrmion number to remain unchanged.
@@
Despite the challenges presented above, we believe that photonic computing using optical Skyrmions remains an exciting avenue to explore. Given the rise of artificial intelligence and machine learning, the need for power-efficient computing technologies is more important than ever. Photonic computing has emerged as a promising solution to meet this need, and optical Skyrmions represent a way of carrying high-dimensional information within optical fields that holds great potential for enhancing information density in photonic computing without additional energy cost. With wavelength, amplitude, and phase information (including structured phase) accessible independent of spatially varying polarization and hence any underlying Skyrmion structure, there are certainly intriguing prospects for combining different existing architectures to exploit all these dimensions simultaneously. Moreover, our work provides a method of directly exploiting the topological and discrete nature of the Skyrmion for computations, and therefore a route to robust digital photonic computing tolerant to perturbations and noise with strong potential for scalability. Most importantly, using optical Skyrmions as units of computation expands the traditional notion of the bit to theoretically infinite values and therefore has the potential to alter the binary foundations of digital computing fundamentally.

### v2

@@
Recent developments in structured light have enabled the generation of optical Skyrmions [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19], which are complex spatially varying polarization fields that carry information through a topological number taking values in the integers. Optical Skyrmions have three crucial properties that make them ideal candidates for high-density data transfer [15], namely the ability to interface with digital information given the discrete nature of the Skyrmion number, a hypothesized robustness to perturbations, and the potential to store arbitrarily large integers within a single localized analog optical field.
@@
These same reasons also make the optical Skyrmion a natural candidate for computation, with its integer-valued topological number offering a route to digital photonic computing in a way that transcends the usual bitwise framework of digital electronics. Moreover, compared to existing photonic computing strategies, which predominantly modulate amplitude and phase, complex spatially varying polarization fields represent an untapped dimension that can be independently manipulated and, therefore, has the potential to increase bandwidth significantly.
@@
The central reason for exploring the use of optical Skyrmions in photonic computing is, however, their hypothesized resistance to noise and perturbations which arises from the integer-valued nature of the Skyrmion number. The topological robustness of Skyrmions has already been established in various domains, including magnetic Skyrmions [20, 21, 22], and more recently, non-local quantum Skyrmions [23, 24], where the idea of using Skyrmions to digitize information is developed from a different perspective. In the setting of optical Skyrmions carried through polarization fields, the study of topological robustness has also recently garnered attention [25], with a degree of protection already verified for propagation through a number of different media [26, 10, 14]. Robustness to noise is particularly relevant to photonic computing, which has thus far remained analog and, therefore, susceptible to noise. Take, for instance, the following three common photonic accelerators. The accumulation of random phase errors in Mach-Zehnder interferometer meshes [27, 28], thermal crosstalk in microring resonator weight banks [29, 30] and low optical contrast in phase-change material-based photonic crossbar arrays [31] all lead to degraded signal-to-noise ratio in large-size implementations, circumventing the scalability of these architectures. Even with progress in device innovation [32, 33], system calibration, and control algorithm optimization to minimize errors[34, 35], almost all demonstrated photonic accelerators thus far have had a size of less than 4 inputs by 4 outputs. Computing using optical Skyrmions is, therefore, a promising solution offering greater scalability through improved stability against noise.
@@
Motivated by this possibility, we describe here a class of structured matter which, when restricted to particular types of Skyrmions, behave as adders/subtractors (Fig. 1) and provide experimental evidence supporting this fact. Note that the same medium can perform both addition and subtraction, but we adopt the term “adder” throughout this paper for brevity. We would also like to emphasize that Skyrmions are not topologically protected through the types of structured matter described in this paper, which have been specially designed to manipulate the Skyrmion number. However, the topological nature of the Skyrmion gives rise to resilience against perturbations in a different form, namely that material parameters have flexibility to fluctuate without affecting the medium’s ability to perform arithmetic, and this resilience of function is of key importance in photonic computing. Such structured matter can be realized using continuously varying retarders with specific structures on their boundary (Extended Fig. 1) and offer a reliable way of manipulating the Skyrmion number without external energy input. Our work opens the doors to this entirely new framework for performing photonic computing, where optical Skyrmions are used as the fundamental unit in computation.
@@
Fig. 1: Concept. An optical Skyrmion of arbitrary order $n$ passing through specially designed passive structured matter can have the effect of addition or subtraction by an arbitrary integer, $n\mapsto n\pm k$. Note that the same medium can be used to perform both addition and subtraction. Such structured matter can be realized in many different ways including metasurfaces, inkjet printing of liquid crystal structures, gradient index systems, direct laser writing of birefringent structures in silica, and more. This figure also depicts the Stokes fields of a standard Néel-type Skyrmion passing through such a medium. Throughout this paper, color is used to represent azimuthal angle on the Poincaré sphere, and saturation to represent height (similar to [15]).
@@ 1 Main
Note our mathematical results also imply that the function of any medium designed in this way depends only on the structure of its boundary and is independent of material properties everywhere else provided they are continuous (excluding certain extreme cases; see Discussion and Methods). This is a reflection of the topological structure of the optical Skyrmion exhibited in matter. From a practical perspective, this suggests a strong robustness of any such adder to physical imperfections of the medium that implements it, which greatly eases fabrication.
To give a concrete example, consider a continuously varying retarder [36], which when restricted to its boundary, is linear and has a retardance of $\pi$. Let $k$ be the number of half-revolutions made by the axis of the retarder traversing counterclockwise along its boundary. Then, one may show
@@
To cascade multiple adders together, one must realign the outer boundary after each operation, which is most easily done using a homogenous waveplate. With our proposed design, this amounts to adding a half waveplate after the medium for addition and a half waveplate before for subtraction, provided we design our components for Skyrmions that are right circularly polarized on their boundary (see Extended Fig. 2a). Therefore, combining an adder with a half waveplate effectively produces a cascadable module that performs addition in one direction and subtraction in the other.
Lastly, to demonstrate the feasibility of our proposed medium, we provide experimentally measured Stokes fields of Skyrmions passing through second-order adders realized via gradient index systems in Fig. 2. We generated Skyrmions of orders ranging from $-3$ to 3 using a cascade of 2 spatial light modulators [37] and passed the field through the medium with appropriate waveplates to achieve the operations $+2+2$, $+2-2$, $-2+2$ and $-2-2$ (see Extended Fig. 2b), where the output field is measured using traditional Stokes polarimetry. Note the medium we used was discarded by quality assurance due to its asymmetric axis distribution, and it exhibits significant perturbations in material properties compared to the typical sample. Despite this, the numerically computed Skyrmion numbers in Fig. 2 show that our proposed adder efficiently and reliably performs the desired operations, even with imperfections in the medium.
@@ 2 Discussion
As mentioned earlier, a crucial property of any such adder is that its effect on Skyrmion fields depends only on the structure of its boundary, a reflection of the topological nature of the optical Skyrmion. From an engineering perspective, this provides clear advantages in manufacturing as the tolerable margins of error in fabrication are greatly relaxed, with no further restrictions on the medium besides the continuity of its material properties. In particular, the function of the adder is independent of perturbations to material properties away from the boundary, including spatially varying anisotropic absorptions with tolerance up to a certain diattenuation and complex spatially varying retardance such as those resulting from birefringence (see [26] for the technical details). We demonstrate the robustness arising from this “topological duality” between field and matter in Extended Fig. 1, which shows the invariance of the effect of our proposed adder on the Skyrmion number of optical fields to disorder in the medium.
@@
Secondly, understanding the non-trivial behavior of polarization fields in waveguides is of key importance for enabling small integrated devices to make use of Skyrmions. In particular, apart from fabricating the adders themselves, it is also necessary to relay polarization information between adders for any meaningful computation to be done. For a waveguide to support the propagation of complex structured fields, it must necessarily be large enough to carry multiple modes. Moreover, work will need to be done to establish the range of Skyrmion numbers that can be achieved by an arbitrary superposition of propagating modes, and the conservation of Skyrmion number in propagation. Here, we provide a heuristic argument suggesting the feasibility of topological protection. Based on the ellipticity of the Helmholtz equation, one has, in general, that the electric field develops continuously within the waveguide. This then naturally descends onto a homotopy of Stokes fields provided there are no zeros of the field. Therefore, at least for some finite distance, we expect the Skyrmion number to remain unchanged.
@@
Despite the challenges presented above, we believe that photonic computing using optical Skyrmions remains an exciting avenue to explore. Given the rise of artificial intelligence and machine learning, the need for power-efficient computing technologies is more important than ever. Photonic computing has emerged as a promising solution to meet this need, and optical Skyrmions represent a way of carrying high-dimensional information within optical fields that holds great potential for enhancing information density in photonic computing without additional energy cost. With wavelength, amplitude, and phase information (including structured phase) accessible independent of spatially varying polarization and, hence, any underlying Skyrmion structure, there are certainly intriguing prospects for combining different existing architectures to exploit all these dimensions simultaneously. Moreover, our work provides a method of directly exploiting the topological and discrete nature of the Skyrmion for computations and, therefore, a route to robust digital photonic computing tolerant to perturbations and noise with strong potential for scalability. Most importantly, using optical Skyrmions as units of computation expands the traditional notion of the bit to theoretically infinite values and, therefore, has the potential to alter the binary foundations of digital computing fundamentally.